OpenStudy (anonymous):

Solve the differential equation xy'(2y - 1) = y(1 - x).

OpenStudy (anonymous):

What happened to your differentials?

OpenStudy (anonymous):

\[\int\limits2dy- \int\limits\frac{dy}{y}=\int\limits\frac{dx}{x}-\int\limits dx +C\]

OpenStudy (anonymous):

Oh I see, break it up. I was trying to use u substitution. So.. 2y - ln[y] = ln[x] -x

OpenStudy (anonymous):

just don't forget the constant C

OpenStudy (anonymous):

Thanks so much! Can I just leave it like that?

OpenStudy (anonymous):

yes